### v3

@@ Abstract
CONCLUSIONS. The ECG foundation model offers an effective solution, allowing it to generalize across a wide range of tasks. By enhancing existing cardiovascular diagnostics and supporting integration into mobile monitoring systems, it contributes to the advancement of the cardiovascular AI community and enables management of cardiac conditions.
@@ 1 Introduction
The trained model and labeled data will be publicly released upon publication through the Brain Data Science Platform (bdsp.io), huggingface and PhysioNet, promoting further research and development in this field.
@@ 2.1 Datasets and pre-processing
Additionally, we utilized the MIMIC-IV-ECG DB to fine-tune our model for various downstream tasks. The MIMIC-IV-ECG dataset is part of the MIMIC series, focusing on the collection and analysis of ECG data.20, 21 MIMIC-IV-ECG originates from real clinical settings at Beth Israel Deaconess Medical Center (BIDMC) in Boston, USA, and contains 800,035 clinical ECGs from 161,352 patients treated in the intensive care unit (ICU). Moreover, ECG recordings in the dataset can be matched with the electronic health records of patients in the MIMIC-ED, allowing ECG data to be associated with specific conditions. Here, we linked several clinical downstream tasks, such as age, sex, chronic kidney disease (CKD), chronic heart disease (CHD), left ventricular ejection fraction (LVEF) and NT-proBNP with ECG data to explore the performance improvement of the fine-tuned ECGFounder model in detecting other diseases and clinical events. More details about the split and labels of MIMIC-IV-ECG DB can be found in Supplement.
@@
For ECG preprocessing, we applied linear interpolation to resample ECG frequencies to 500Hz. We used a high-pass filter with a cutoff frequency of 1 Hz to suppress residual baseline drift and a second-order 30 Hz Butterworth low-pass filter to reduce high-frequency noise. A 50/60 Hz notch filter was utilized to eliminate electrical interference. From ECG records longer than 10 seconds, we extracted 10-second windows in sequence. If a sequence was less than 10 seconds, we applied zero padding. We normalized all signals using the mean and standard deviation of each individual signal segment before inputting them into the model. If a signal was missing, we assigned it a value of 0.
@@ 2.2 Model architecture
To establish an ECG foundation model, we require an architecture that is tailored for ECG, capable of learning generalizable representations from large-scale ECG datasets. The growth in the size of ECG data and the increase in the number of leads mean that the model must not only consider temporal information of the data but also spatial relationships, i.e., interactions between different leads and the overall pattern of cardiac electrical activity. This is crucial for the ECG foundation model to be applicable in real clinical ECG scenarios, to mitigate the impact of non-uniform ECG durations and missing leads in the training dataset.
Considering these factors, we built our model architecture on top of the RegNet architecture.23, 24 This structure begins with a stage-wise network design where each stage consists of a set number of blocks and channels that scale with network depth. This is beneficial for us to expand and design the blocks and channels of the ECG foundation model. Unlike traditional uniform scaling across layers, RegNet employs a quantized linear model to predict optimal widths and depths, ensuring efficient performance across a range of model sizes. The initial layers focus on capturing low-level features with fewer channels, which gradually increase as the network deepens, thus optimizing computational efficiency and model capacity. Following this, the model utilizes a series of bottleneck blocks that combine group convolutions and channel-wise attention mechanisms, enhancing the richness of representation in both temporal and spatial dimensions while controlling model complexity. This tailored configuration makes the model suitable for ECG data. More details about the model architecture can be found in the Supplement.
@@ 2.3 Training with real-world annotations: noisy, imbalanced, positive unlabeled
Unlike conventional ECG diagnostic models that typically use single-label classification methods, we employed multi-label classification during the training phase of the ECG foundation model. This approach aligns more closely with clinical practice, where an ECG diagnosis often consists of multiple diagnostic labels. For example, an abnormal ECG diagnosis might be something like normal sinus rhythm | premature ventricular complexes | premature ectopic complexes. Additionally, multi-label diagnostics better meet the application needs of clinicians. During training, utilizing a multi-label classification approach provides our model with rich semantic value and facilitates generalization to different annotation systems. However, the nature of multi-label annotations means that it is almost impossible to achieve perfect multi-label data, especially in cases like ECG where there are many diagnostic categories. Generally speaking, if a cardiology expert can annotate three classes simultaneously, it is considered excellent. However, the actual number of ECG diagnostic labels far exceeds three classes.25 To address the challenge brought by incomplete ECG annotations, we introduce positive unlabeled (PU) learning method.
PU learning is defined as when positive samples are present in the data but unlabeled, but the labeled positive samples are correct, meaning that the unlabeled samples are not necessarily negative examples.26 Typically, PU learning leads to a severe imbalance in the predicted probability distribution: an ECG usually contains a few positive diagnostic labels, with the remaining labels being treated as negative, leading to a severe positive-negative imbalance where the negative samples far outnumber the positive samples for each label. The inherent scarcity of positive examples, combined with the missing labels, exacerbates the imbalance, making it more challenging to learn from the positive examples. When using conventional loss functions, such as the Binary Cross-Entropy (BCE) loss function, the model tends to learn from simpler samples, namely the true negative (TN) samples, while the more challenging samples, namely the false negative (FN) samples, are harder to fit. When the model is biased towards learning TN samples, its predicted probabilities tend to skew towards 0 rather than 1, which diminishes its ability to predict positive samples and weakens its potential to detect clinically important abnormalities.
@@ 2.4 Training a single-lead ECG model based on leads augmentation
The development of portable and wearable ECG devices in recent years has revolutionized continuous cardiac monitoring, offering a non-invasive method for real-time health assessment. Beyond the conventional diagnosis of cardiac arrhythmias, another critical challenge in this field is accurately detecting and interpreting ECG axis deviation on single-lead ECGs from wearable devices (typically lead I), which can significantly impact the diagnosis of various cardiac abnormalities. For instance, left axis deviation can provide additional insights into diagnoses like left ventricular hypertrophy, left bundle branch block, left anterior fascicular block, pre-excitation syndromes, and inferior myocardial infarction.27 To address this issue, we developed a novel training method utilizing lead-augmented wearable ECG models. By systematically enhancing standard 12-lead ECG data, we simulate various clinical scenarios of axis inversion, thereby enhancing the model’s robustness and versatility. Understanding the relationship between ECG vectors and leads is crucial for this method. The cardiac electrical activity generates a vector representation of cardiac signals, captured by different ECG leads placed on the body. Each lead provides a unique perspective of the cardiac electrical axis, offering a comprehensive view when combined. The standard 12-lead ECG system includes limb leads (I, II, III, aVR, aVL, aVF) and precordial leads (V1-V6), with each lead representing a specific projection of the cardiac electrical vector.28 Specifically, we primarily utilize the ECG signals from limb leads. Based on the angular position of each limb lead’s axis relative to the heart, we consider lead I as the center of the semicircle (i.e., 0°) and calculate the signals from all leads around the semicircle (i.e., from -90° to 90°), thereby obtaining six additional leads (aVL, -aVR, II, -III, aVF, -aVF). We then trained a model for wearable ECG devices, extracting the lead I ECG from the HEEDB 12-lead data and randomly incorporating one of the remaining six augmented leads into the model for training with a 50% probability. This ensures that the model can learn arrhythmia features from lead I data and axis deviation from the additional six leads. Additionally, we have scaled down the model’s parameter size to optimize for wearable devices with limited computational resources.
@@ 2.5 Fine-tuning on ECG foundation model
When adapting to specific ECG downstream tasks, we need to retain the parameters of the base model and discard the initial classification linear layer. The number of classes in the downstream task determines the number of neurons needed in the final layer of the new linear layer. The training objective is to generate classification outputs that match the labels.
@@ 2.6 Clinical validation
To validate and compare the performance of our model, we established a committee consisting of three experienced ECG physicians to annotate the internal test set. We developed an ECG annotation system for physicians to use, and set up 20 types of labels including cardiac rate abnormalities, conduction blocks, myocardial dilation, myocardial infarction, and ECG morphologies, with several sub-labels under each category. The Supplement displays the complete list of label types. After independent annotation by the committee, a consensus determination was made; ECGs that did not reach consensus were removed, providing an expert standard for model evaluation. ECGs that the committee could not interpret or agree upon were eliminated from our test dataset.
Additionally, to compare the diagnostic accuracy between the model and experts, five additional ECG physicians were involved in this study. They were given specific instructions on how to use the system. The Supplement shows the doctors’ ages, levels of experience, and education. Each doctor was required to independently annotate each ECG from the internal test set. The annotations from these physicians were then compared to the model’s results.
The evaluation of the model was conducted by calculating accuracy, the area under the Receiver Operating Characteristic curve (AUROC), sensitivity, and specificity, each with bilateral 95% confidence intervals (CI).
@@ 3 Results
Our dataset comprises 10,771,552 expert-annotated ECGs recorded from 1,818,247 unique subjects using 12-lead ECG machines, covering the period from the 1990s to present, with a recording duration of 10 seconds. The average age of patients is 60.6 years, with 48.7% being female.15 The development dataset includes 7,519,035 ECGs from 1,725,643 patients, and the test dataset includes 834,926 ECGs from 265,818 patients. The committee internal test set includes 523 ECGs recorded from 2023, corresponding to 523 unique patients. We obtained a total of 150 labels describing the ECGs, including heart rate abnormalities, conduction abnormalities, cardiac hypertrophy, myocardial infarction, myocardial ischemia, pacemaker devices, and ECG waveform morphology. Using these labels, we trained two deep learning ECG models separately: one for standard 12-lead ECGs and the other designed for single-lead wearable device ECGs (lead I). We evaluated the performance of both models on internal and external test sets.
We first conducted large-scale validation of ECGFounder using an internal test dataset comprising 834,926 ECGs, to assess its consistency with diagnoses confirmed by clinically recognized cardiologists (Figure S3). ECGFounder demonstrated an impressive AUC of at least 0.950 across 82 diagnostic categories and at least 0.900 across 112 categories. Overall, the AUCs were particularly high for categories such as rhythm disorders, conduction abnormalities, cardiac hypertrophy, and myocardial infarction. In contrast, the AUCs were lower for categories involving waveform changes, which may be attributed to the annotation habits of certain physicians who tend to place less emphasis on documenting waveform changes in their diagnoses.
@@
We validated the deep learning model for 12-lead ECGs using the committee internal test set. The algorithm’s average AUC score for diagnosing all 20 classifications was 0.968 (95% CI, 0.965–0.969), sensitivity was 0.971 (95% CI, 0.970–0.973), and specificity was 0.937 (95% CI, 0.935–0.938). When comparing the model with cardiologists, our model achieved an overall average F1 score of 0.677, outperforming the cardiologists’ average F1 score of 0.640. The model’s performance was compared with cardiologists’ performance across 20 diagnostic categories (Table S7 in Supplement). When comparing the model’s ROC curve with the true positive rate and false positive rate of cardiologists, the model outperformed the average performance of cardiologists for most labels (Figure S2 in Supplement).
@@
In the external test set experiments for 12-lead ECGs, we evaluated the performance of our model and other models on the CODE-test and PTB-XL datasets. On CODE-test, our model achieved an average AUC of 0.981, outperforming other baseline models such as CinC2020-CTN, which had an average AUC of 0.963, and 0.980 (S12L-ECG) (Table 2). On PTB-XL, because the other two models can only diagnose arrhythmias and cannot complete other diagnostic classifications in the broader classes on PhysioNet Challenge-2020, we validated this dataset only on the class which the models have. Our model achieved an average F1 score of 0.732 and an average AUC of 0.924 (Table 2). These results indicate that our model generalizes to different regions, hospitals, and patients.
@@
In the internal test set experiments focused on single-lead ECGs, the model demonstrated excellent performance in rhythm-type diagnoses(Figure S3). Notably, it achieved an AUC above 0.95 for common heart rate abnormalities such as normal sinus rhythm, sinus bradycardia, sinus tachycardia, marked sinus bradycardia, sinus arrhythmia, marked sinus arrhythmia, and atrial fibrillation. These diagnoses can be reliably identified from single-lead ECGs. Additionally, the model achieved an AUC above 0.8 for diagnoses including premature ventricular complexes, premature supraventricular complexes, pacemaker, 1st degree AV block, branch block, fascicular block, and chamber enlargement. It is worth noting that for some of these diagnoses, this is the first time such exceptional performance has been observed using single-lead ECGs. In the diagnosis of myocardial infarctions, the model also showed good diagnostic performance for lateral infarct, anterolateral infarct, acute MI, and STEMI. In ECG research, these diagnoses are associated with shifts in the heart’s electrical axis. This suggests that data augmentation methods based on the electrical axis are effective for training single-lead ECG models.
@@
In the external test set experiments on single-lead ECG devices, our model accurately classified sinus rhythm and atrial fibrillation. The model’s performance, as shown in the table, achieved an AUC of 0.975 and 0.957 for normal sinus rhythm and atrial fibrillation, respectively (Table S8 in Supplement). This indicates that the model achieves excellent performance on portable and wearable ECGs in real-world scenarios.
@@
We next validated the performance of ECGFounder in transfer learning. Our model was fine-tuned and evaluated on six downstream tasks using supervised learning on the MIMIC-IV-ECG dataset, resulting in both a 12-lead ECG model and a single-lead ECG model. The fine-tuning results are shown in Figures 2. As shown, ECGFounder outperforms the baseline methods in every downstream task. Specifically, it achieves 2 to 3 percentage points higher performance than ECG-SimCLR and 4 to 6 points higher than ECG-ResNet in age, sex, NTproBNP, CKD, and CHD detection. Additionally, to compare with the previously published CKD study 29, we created an independent test set consisting of individual patients from the MIMIC-IV-ECG dataset for validation. As shown in the results (Table 3), our model surpasses the previously studied model in all six classifications for both 12-lead and single-lead ECGs.

### v4

@@ Abstract
CONCLUSIONS. The ECG foundation model offers an effective solution, allowing it to generalize across a wide range of tasks. By enhancing existing cardiovascular diagnostics and facilitating integration with cloud-based systems that analyze ECG data uploaded from wearable devices, it significantly contributes to the advancement of cardiovascular AI community and enables management of cardiac conditions. Code and model can be found at https://github.com/PKUDigitalHealth/ECGFounder.
@@ 2.1 Datasets and pre-processing
Additionally, we utilized the MIMIC-IV-ECG DB to fine-tune our model for various downstream tasks. The MIMIC-IV-ECG dataset is part of the MIMIC series, focusing on the collection and analysis of ECG data.20, 21 MIMIC-IV-ECG originates from real clinical settings at Beth Israel Deaconess Medical Center (BIDMC) in Boston, USA, and contains 800,035 clinical ECGs from 161,352 patients treated in the intensive care unit (ICU). Moreover, ECG recordings in the dataset can be matched with the electronic health records of patients in the MIMIC-ED, allowing ECG data to be associated with specific conditions. Here, we linked several clinical downstream tasks, such as age, sex, chronic kidney disease (CKD), chronic heart disease (CHD), left ventricular ejection fraction (LVEF) and NT-proBNP with ECG data to explore the performance improvement of the fine-tuned ECGFounder model in detecting other diseases and clinical events. More details about the split and labels of MIMIC-IV-ECG DB can be found in Supplement. Proportions of positive and negative cases naturally reflect clinical prevalence but were not deliberately controlled or balanced.
@@
For data preprocessing, unreadable files, missing data, and unmatched data were excluded. Our final development dataset includes 7,519,035 ECGs from 1,319,128 patients, and the test dataset includes 834,926 ECGs from 146,570 patients. We applied linear interpolation to resample ECG frequencies to 500 Hz. We used a high-pass filter with a cutoff frequency of 0.5 Hz to suppress residual baseline drift and a second-order 50 Hz Butterworth low-pass filter to reduce high-frequency noise. A 50/60-Hz notch filter was utilized to eliminate electrical interference. For ECG records longer than 10 seconds, we extracted 10-second windows in sequence. If a sequence was less than 10 seconds, we applied zero padding. We normalized all signals using the mean and standard deviation of each individual signal segment before inputting them into the model.
@@ 2.2 Model architecture
To establish the model, we used an architecture tailored for ECG, capable of learning generalizable representations from large-scale ECG datasets. The increase of ECG data and the number of leads meant that the model must not only consider temporal information, but also spatial relationships (i.e., interactions between different leads and the overall pattern of cardiac electrical activity). This is crucial for ensuring that the ECG foundation model is applicable to real-world clinical ECG scenarios, as it mitigates the impact of nonuniform ECG durations and missing leads in the training dataset.
Considering these factors, we built our model architecture based on our previously proposed Net1D.23 It is built on top of the Self-Regulated Network for Image Classification (RegNet) architecture.24 This structure begins with a stage-wise network design where each stage consists of a set number of blocks and channels that scale with network depth. This is beneficial for us to expand and design the blocks and channels of the ECG foundation model. Unlike traditional uniform scaling across layers, RegNet employs a quantized linear model to predict optimal widths and depths, ensuring efficient performance across a range of model sizes. The initial layers focus on capturing low-level features with fewer channels, which gradually increase as the network deepens, thus optimizing computational efficiency and model capacity. Following this, the model utilizes a series of bottleneck blocks that combine group convolutions and channel-wise attention mechanisms, enhancing the richness of representation in both temporal and spatial dimensions while controlling model complexity. This tailored configuration makes the model suitable for ECG data. More details about the model architecture can be found in the Supplement.
@@ 2.3 Training with real-world annotations: noisy, imbalanced, positive unlabeled
Unlike conventional ECG diagnostic models that typically use single-label classification methods, we employed multi-label classification during the training phase of the ECG foundation model. This approach aligns more closely with clinical practice, where an ECG diagnosis often consists of multiple diagnostic labels. For example, an abnormal ECG diagnosis might be something like normal sinus rhythm | premature ventricular complexes | premature ectopic complexes. Additionally, multi-label diagnostics better meet the application needs of clinicians. During training, utilizing a multi-label classification approach provides our model with rich semantic value and facilitates generalization to different annotation systems.
However, the nature of multi-label annotations means that it is almost impossible to achieve perfect multi-label data, especially in cases like ECG where there are many diagnostic categories. Generally speaking, if a cardiology expert can annotate three classes simultaneously, it is considered excellent. However, the actual number of ECG diagnostic labels far exceeds three classes.25
To address the challenge brought by incomplete ECG annotations, we introduce positive unlabeled (PU) learning method. PU learning is defined as when positive samples are present in the data but unlabeled, but the labeled positive samples are correct, meaning that the unlabeled samples are not necessarily negative examples.26 Conventional multilabel classification methods typically treat unlabeled categories as negative by default, but this assumption does not hold true for ECG data.
Positive unlabeled learning in the ECG context leads to a severe imbalance in the predicted probability distribution: an ECG usually contains a few positive diagnostic labels, with the remaining labels being treated as negative. This creates a severe positive–negative imbalance, where negative samples far outnumber positive samples for each label. When using conventional loss functions, such as the binary cross-entropy loss function, the model tends to learn from simpler samples, namely the true negative samples, while the more challenging samples, namely the false negative samples (which may in fact be missed positives), are harder to fit. As a result, the model’s predicted probabilities tend to skew toward 0 rather than 1, diminishing its ability to detect clinically important abnormalities.
@@ 2.4 Training a single-lead ECG model based on leads augmentation
In recent years, portable and wearable ECG-device development has revolutionized continuous cardiac monitoring, offering a noninvasive method for real-time assessment. Beyond the conventional diagnosis of cardiac arrhythmias, another critical challenge in this field is accurately detecting and interpreting ECG axis deviation on single-lead ECGs from wearable devices (typically lead I), which can significantly impact the diagnosis of various cardiac abnormalities. For instance, left axis deviation can provide additional insights into diagnosis, such as left ventricular hypertrophy, left bundle-branch block, left anterior fascicular block, preexcitation syndromes, and inferior myocardial infarction (MI).27
To address this issue, we developed a novel training method utilizing lead-augmented wearable ECG models. By systematically enhancing standard 12-lead ECG data, we simulate various clinical scenarios of axis inversion, thereby enhancing the model’s robustness and versatility. Understanding the relationship between ECG vectors and leads is crucial for this method. The cardiac electrical activity generates a vector representation of cardiac signals, captured by different ECG leads placed on the body. Each lead provides a unique perspective of the cardiac electrical axis, offering a comprehensive view when combined. The standard 12-lead ECG system includes limb leads (I, II, III, aVR, aVL, aVF) and precordial leads (V1-V6), with each lead representing a specific projection of the cardiac electrical vector.28 Specifically, we primarily utilize the ECG signals from limb leads. Based on the angular position of each limb lead’s axis relative to the heart, we consider lead I as the center of the semicircle (i.e., 0°) and calculate the signals from all leads around the semicircle (i.e., from -90° to 90°), thereby obtaining six additional leads (aVL, -aVR, II, -III, aVF, -aVF). We then trained a model for wearable ECG devices, extracting the lead I ECG from the HEEDB 12-lead data and randomly incorporating one of the remaining six augmented leads into the model for training with a 50% probability. This ensures that the model can learn arrhythmia features from lead I data and axis deviation from the additional six leads. Additionally, we have scaled down the model’s parameter size to optimize for wearable devices with limited computational resources.
@@ 2.5 Fine-tuning on ECG foundation model
When adapting to specific ECG downstream tasks, we needed to retain the parameters of the base model and discard the initial classification linear layer. The number of classes in the downstream task determines the number of neurons needed in the final layer of the new linear layer. The training objective is to generate classification outputs that match the labels. We adopted two different methods of fine-tuning: linear probing and full fine-tuning. During the linear probing experiments, we only fine-tuned the parameters of the linear classification head on top of the pretrained model, keeping all other pretrained model weights frozen. During full fine-tuning, we allowed all pretrained model weights to be updated and adapted to the downstream classification tasks.
@@ 2.6 Clinical validation
To validate and compare the performance of our model, we followed the committee experimental design of Hannun et al.29 We established the committee consisting of three experienced ECG cardiologists to annotate a subset of the internal test set, which includes the 523 most recent ECGs from 523 unique patients. We developed an ECG annotation system for cardiologists and 20 label types, including cardiac-rate abnormalities, conduction blocks, myocardial dilation, MI, and ECG morphologies, with sublabels under each category. Table 1 displays the complete list of label types. After independent annotation by the committee, a consensus determination was made; labels that did not reach consensus were removed, providing an expert standard for model evaluation. Labels that the committee could not interpret or agree on were eliminated from our test dataset.
In addition, to compare the diagnostic accuracy between the model and cardiologists, four additional ECG cardiologists were involved and provided specific instructions on how to use the system. Table S6 shows the cardiologists’ ages, levels of experience, and education. Each cardiologist was required to annotate each ECG from the previous internal test set. The annotations from these cardiologists were then compared with the model’s results.
The evaluation of the model was conducted by calculating accuracy, the area under the Receiver Operating Characteristic curve (AUROC), sensitivity, and specificity.
@@ 3 Results
Experimental results demonstrate that ECGFounder achieves superior performance on internal validation sets of 12-lead ECGs, with AUROC exceeding 0.95 for 80 diagnoses (Fig. S3). We further validated ECGFounder for 12-lead ECGs using the committee’s internal test set. The algorithm’s average AUROC score for diagnosing all 20 classifications was 0.968 (95% confidence interval [CI], 0.955 to 0.982), sensitivity was 0.971 (95% CI, 0.639 to 0.988), and specificity was 0.937 (95% CI, 0.912 to 0.953). When comparing the model with cardiologists, our model achieved an overall average F1 score of 0.677 (95% CI, 0.480 to 0.802), outperforming the cardiologists’ average F1 score of 0.640. The model’s performance was compared with cardiologists’ performance across 20 diagnostic categories (Table 1). When comparing the model’s receiver operating characteristic curve with the true-positive rate and false-positive rate of cardiologists, the model outperformed the average performance of cardiologists for most labels (Fig. S2).
@@
In the external test set experiments for 12-lead ECGs, we evaluated the performance of our model and other models on the CODE-test and PTB-XL datasets. On CODE-test, our model achieved an average AUROC of 0.981 (95% CI, 0.979 to 0.984), outperforming other baseline models such as S12L-ECG, which had an average AUROC of 0.980 (95% CI, 0.978 to 0.982); CTN, which had an average AUROC of 0.963 (95% CI, 0.960 to 0.967); and ECG Squeeze-and-Excitation Residual Neural Network (ECG-SE-ResNet), which had an average AUROC of 0.963 (95% CI, 0.961 to 0.967) (Table 2). On PTB-XL, as the other two models can only diagnose arrhythmias and cannot complete other diagnostic classifications, we validated this dataset only on the class that the models have. Our model achieved an average AUROC of 0.924 (95% CI, 0.917 to 0.931) (Table 2).17, 30, 31 These results indicate that our model generalizes to different regions, hospitals, and patients.
@@
Internal test set experiments focused on single-lead ECGs, with the model demonstrating excellent performance in rhythm-type diagnosis (Fig. S3). It achieved an AUROC above 0.95 for common heart-rate abnormalities such as normal sinus rhythm, sinus bradycardia, sinus tachycardia, marked sinus bradycardia, sinus arrhythmia, marked sinus arrhythmia, and AF. These diagnoses can be reliably identified from single-lead ECGs. In addition, the model achieved an AUROC above 0.8 for diagnosis, including premature ventricular complexes, premature supraventricular complexes, pacemaker, first-degree atrioventricular block, branch block, fascicular block, and chamber enlargement. The performance observed for some of these diagnoses was exceptional for the use of single-lead ECGs and is worthy of note. In MI diagnosis, the model also showed good diagnostic performance for lateral infarct, anterolateral infarct, acute MI, and ST-segment elevation MI. In ECG research, these diagnoses are associated with shifts in the heart’s electrical axis. This suggests that data-augmentation methods based on the electrical axis are effective for training single-lead ECG models.
@@
In the external test set experiments on single-lead ECG devices, our model accurately classified sinus rhythm and AF. The model’s performance, as shown in supplement, achieved an AUROC of 0.975 (95% CI, 0.972 to 0.977) and 0.957 (95% CI, 0.955 to 0.959) for normal sinus rhythm and AF, respectively. These results demonstrate excellent performance in analyzing ECG signals collected from portable and wearable ECG devices under real-world conditions. It should be emphasized that ECGFounder is executed on a cloud-based system, analyzing data uploaded from wearable devices, rather than operating directly on the wearable devices themselves.
We next validated the performance of ECGFounder in transfer learning. Our model was fine-tuned and evaluated on six downstream tasks using supervised learning on the MIMIC-IV-ECG dataset, resulting in both a 12-lead ECG model and a single-lead ECG model. The fine-tuning results are shown in Figure 2. As shown, ECGFounder outperforms the baseline methods in every downstream task. Specifically, it achieves 2 to 3 percentage points higher performance than ECG-SimCLR and 4 to 6 percentage points higher than ECG- in age, sex, NT-proBNP, LVEF, CKD, and CHD detection. In addition, for comparison with the previously published CKD study,32 we created an independent test set consisting of individual patients from the MIMIC-IV-ECG dataset for validation. The results are shown in Table 3. It should be noted that ECGFounder was the internal validation, while the compared method was the external validation.
@@
We next validated the performance of ECGFounder in transfer learning. Our model was fine-tuned and evaluated on six downstream tasks using supervised learning on the MIMIC-IV-ECG dataset, resulting in both a 12-lead ECG model and a single-lead ECG model. The fine-tuning results are shown in Figures 2. As shown, ECGFounder outperforms the baseline methods in every downstream task. Specifically, it achieves 2 to 3 percentage points higher performance than ECG-SimCLR and 4 to 6 points higher than ECG-ResNet in age, sex, NTproBNP, CKD, and CHD detection. Additionally, to compare with the previously published CKD study 32, we created an independent test set consisting of individual patients from the MIMIC-IV-ECG dataset for validation. As shown in the results (Table 3), our model surpasses the previously studied model in all six classifications for both 12-lead and single-lead ECGs.